OpenStudy (anonymous):

Describe the differences between the graph of y = –3(x + 7)^2 – 10 and the standard position graph of y = x^2.

OpenStudy (anonymous):

Nevermind, I figured it out. (:

OpenStudy (anonymous):

y=–3(x+7)^2–10 is in General Form, while y=x^2 is Simplest Form. When finding the points to plot for the one in General Form I got the points: (-9,-22), (-8,-13), (-7,-10) Vertex, (-6,-13), and (-5,-22). This graph obviously opens downward, making the vertex at (-7,-10) the maximum (or highest) point. I then substituted the values -2, -1, 0, 1, 2 into the Simplest Form equation. The points I got were: (2,4), (1,1), (0,0) Vertex, (-1,1), (-2,4). This graph opens upward with the vertex (0,0) being the lowest point. Therefore, the difference is that y=–3(x+7)^2–10 opens downward with the vertex being the highest point and y=x^2 opens upward with the vertex being the lowest point.

OpenStudy (anonymous):

Sound right?

OpenStudy (anonymous):

Its FINE

OpenStudy (anonymous):

Fine as in correct?

OpenStudy (yrelhan4):

looks good.

OpenStudy (anonymous):

Thank you